# A Simple, Optimal and Efficient Algorithm for Online Exp-Concave Optimization

### v1

@@ 3.2 The Improper-to-Proper Conversion
LightONSpreserves the infrequent-projection property of LightONS.Core, yielding a total runtime better than both ONS and OQNS. More importantly, LightONS retains the flexible mirror-descent structure of ONS. In fact, LightONS differs from ONS in two key aspects: (i) the projection-hysteresis mechanism (introduced in LightONS.Core to enhance efficiency) and (ii) the domain conversion (introduced here to ensure proper learning), both of which are largely orthogonal to the mirror-descent update in the ONS algorithm. Consequently, LightONS applies to a wide range of scenarios where ONS is essential (particularly when its mirror-descent update plays a critical role), such as gradient-norm adaptive regret, parametric stochastic bandits, and memory-efficient online learning. We illustrate these applications in Section˜5.
@@
3.3 Numerical Implementation
The overall efficiency of LightONS hinges on two key operations: the matrix inversion $A_{t}^{-1}$ and the (infrequent) Mahalanobis projection $\Pi^{A_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{y}}_{t+1}]$. In this subsection, we detail efficient numerical approaches for the two operations respectively.
@@ 3.3 Numerical Implementation
In LightONS, all Mahalanobis projections are onto the Euclidean ball $\mathcal{B}(D/2)$ rather than the potentially complex domain $\mathcal{X}$. This geometry enables customized numerical approaches faster than generic solvers (Lee et al., 2015; Jiang et al., 2020). Specifically, the dual problem of the Mahalanobis projection reduces to a one-dimensional root-finding problem (Exercise 4.22 of Boyd and Vandenberghe (2004)), solvable via bisection. We propose FastProj in Algorithm˜4, with its theoretical guarantees given in the lemma below. The proof of this theorem is provided in Section˜B.5.
Theorem 3.
Let $\mathbf{v}$ denote the output of FastProj (Algorithm˜4), and $\mathbf{v}^{*}=\Pi^{A}_{\mathcal{B}(R)}[\mathbf{u}]$ denote the exact Mahalanobis projection, then $\mathbf{v}\in\mathcal{B}(R)$ and $\left\|\mathbf{v}-\mathbf{v}^{*}\right\|_{2}\leq\zeta$.
With choice 1, the total runtime is $O(d^{\omega}n)$, and with choice 2, the total runtime is $O(d^{3}+dn)$. The number of bisection iterations $n=O(\log(\frac{1}{\zeta}\frac{~\overline{\lambda}~}{~\underline{\lambda}~}))$. In the context of LightONS (Algorithm˜3), $\zeta_{t}=O(1/t^{2})$, $\underline{\lambda}_{t}=\epsilon$ and $\overline{\lambda}_{t}=\epsilon+c_{g}^{2}G^{2}t$, thus $n_{t}=O(\log t)$.
@@
We present two choices for different purposes. Choice 1 attains superior theoretical dependence on $d$ by exploiting fast matrix-multiplication with exponent $\omega<2.3714$. Choice 2 offers stronger practical performance via tridiagonalization (Parlett, 1998; Golub and Van Loan, 2013), which is preferable when adopting the practical setting $\omega=3$.
@@
The accumulated truncation error contributes only a negligible additive constant to the regret of LightONS. Full details are deferred to Section˜B.4.
(a) Instantaneous regret.
(b) Regret.
(c) Norm.
(d) Instantaneous regret.
(e) Regret.
(f) Norm.
Figure 1: Experimental results with $\mathcal{X}=\mathcal{B}(1)$, $d=10$, $T=10^{4}$. The first row shows linear regression with $G=\frac{1}{10}$, $\alpha=5$; The second row shows logistic regression with $G=\frac{1}{10}$, $\alpha=\mathrm{e}^{-1/5}$.
3.4 Empirical Validation
We conduct numerical experiments to validate the theoretical guarantees of LightONS, especially its negligible statistical gap from ONS and its non-asymptotic statistical advantages over OQNS. Three algorithms, namely LightONS, ONS, and OQNS, are configured with their theoretically optimal parameters. The implementation of ONS and LightONS follows Algorithms˜1 and 3, respectively; While OQNS is implemented in accordance with Algorithm 3 of Mhammedi and Gatmiry (2023). We evaluate on two fundamental tasks: linear regression and logistic regression. The loss functions are defined as:
$\ell^{\text{linear}}_{t}(\mathbf{w})=\frac{1}{2}\left(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}+y_{t}\right)^{2},\quad\ell^{\text{logistic}}_{t}(\mathbf{w})=\log\left(1+\exp\left(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}\right)\right),\quad(\mathbf{x}_{t},y_{t})\sim\left|\mathcal{N}(\bm{0},I)\right|.$
Our logistic regression setup slightly differs from the standard formulation, as we omit the binary label $y_{t}\in\left\{0,1\right\}$ for simplicity, which does not affect exp-concavity. Each entry of $\mathbf{x}_{t}$ and $y_{t}$ are i.i.d. sampled from a folded Gaussian distribution, i.e., the absolute value of a standard Gaussian random variable. All results are averaged over 5 independent runs with the same seeds. In Figure˜1, the averaged performance is shown as a dark line, while individual runs are in transparent lines.
Our experiments confirm that computational gains of LightONS incur negligible loss in statistical performance. As shown in regret plots, curves of LightONS (red dotted) overlap with those of ONS (green solid). This provides strong empirical evidence that LightONS retains the sharp regret constants of the ONS. In contrast, OQNS (blue dashed) shows noticeably higher regret and slower convergence in both regret and decision-norm plots, indicating that its theoretical bounds may carry greater constants.
The key advantage of LightONS, its computational efficiency, is evident in decisions-norm plots. After an initial convergence phase (e.g., in Figure˜1(f), the first $10^{2}$ rounds), both ONS and LightONS stabilize near the optimal solution. Then ONS frequently pushes decisions outside the domain, triggering costly Mahalanobis projections. In contrast, LightONS consistently outputs decisions with $\left\|\mathbf{y}_{t}\right\|\leq kD/2=2$ and avoids Mahalanobis projections.
4 Answering a COLT’13 Open Problem
In this section, we highlight the broader impact of our method by extending it to a related setting. Via the online-to-batch conversion, our method applies to Stochastic eXp-concave Optimization (SXO), where the optimal OXO regret translates into the (near) optimal SXO sample complexity while substantially reducing the computational cost. This extension answers a COLT’13 open problem posed by Koren (2013), demonstrating our method’s significance beyond OXO.
4.1 Restatement of the Open Problem
SXO seeks to minimize an exp-concave function $F:\mathcal{X}\to\mathbb{R}$ over a convex domain $\mathcal{X}\subseteq\mathbb{R}^{d}$, where the learner has access to $F$ only through some stochastic oracle. In this section, we consider the stochastic gradient oracle, which returns unbiased estimates of the gradient of $F$, and we are interested in the sample complexity and total runtime required to find an $\epsilon$-optimal solution $\mathbf{x}_{\epsilon}$, i.e.,
$F(\mathbf{x}_{\epsilon})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\leq\epsilon,\quad\text{where}~F(\mathbf{x})=\mathop{\mathbb{E}}_{\xi\sim\mathcal{D}_{\xi}}\left[f(\mathbf{x};\xi)\right].$
The regularity of the stochastic functions is formally stated below (Koren, 2013).
Assumption 4.
The loss function $F:\mathcal{X}\to\mathbb{R}$ is the expectation of a random function $f:\mathcal{X}\times\Xi\to\mathbb{R}$ over an unknown distribution $\mathcal{D}_{\xi}$ on $\Xi$, i.e., $F(\mathbf{x})=\mathop{\mathbb{E}}_{\xi\sim\mathcal{D}_{\xi}}\left[f(\mathbf{x};\xi)\right]$. For any $\xi\in\Xi$, the stochastic loss function $f(\cdot;\xi)$ is $\alpha$-exp-concave, differentiable and $G$-Lipschitz over $\mathcal{X}$. For any query point $\mathbf{x}\in\mathcal{X}$, the stochastic gradient oracle returns $\nabla f(\mathbf{x};\xi)$ with $\xi$ i.i.d. drawn from $\mathcal{D}_{\xi}$.

### v2

@@ 3.2 The Improper-to-Proper Conversion
LightONSachieves a total runtime better than both ONS and OQNS while matching the regret of ONS. More importantly, LightONS retains the flexible OMD framework of ONS. In fact, LightONS differs from ONS in two key aspects: (i) the deferred-projection mechanism (introduced in LightONS.Core to boost efficiency) and (ii) the domain conversion (introduced here to ensure proper learning), both of which are largely orthogonal to the mirror-descent update in the standard ONS. Consequently, LightONS applies to various scenarios where ONS is essential, particularly when its mirror-descent update plays a critical role. We illustrate these applications in Section˜5.
4 Answering a COLT’13 Open Problem
Via the online-to-batch conversion, our method applies to Stochastic eXp-concave Optimization (SXO), where the optimal OXO regret translates into the (near) optimal SXO sample complexity while substantially reducing the computational cost. This extension answers a COLT’13 open problem posed by Koren [2013], demonstrating our method’s significance beyond OXO.
4.1 Restatement of the Open Problem
SXO seeks to minimize an exp-concave function $F:\mathcal{X}\to\mathbb{R}$ over a convex domain $\mathcal{X}\subseteq\mathbb{R}^{d}$, and the learner has access to $F$ only through some stochastic oracle. In this section, we consider the case where $F$ is the expectation of a random function $f:\mathcal{X}\times\Xi\to\mathbb{R}$ over a distribution on $\Xi$, and the learner has access to the gradient of $f(\cdot;\xi)$ with $\xi$ drawn from $\mathcal{D}_{\xi}$. We are interested in the sample complexity and total runtime required to find an $\epsilon$-optimal solution $\mathbf{x}_{\epsilon}$, i.e.,
$F(\mathbf{x}_{\epsilon})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\leq\epsilon,\quad\text{where}\penalty 10000\ F(\mathbf{x})=\mathop{\mathbb{E}}_{\xi\sim\mathcal{D}_{\xi}}\left[f(\mathbf{x};\xi)\right].$
The regularity of the random function is formally stated below.
Assumption 4.
The loss function $F:\mathcal{X}\to\mathbb{R}$ is the expectation of a random function $f:\mathcal{X}\times\Xi\to\mathbb{R}$ over an unknown distribution $\mathcal{D}_{\xi}$ on $\Xi$, i.e., $F(\mathbf{x})=\mathop{\mathbb{E}}_{\xi\sim\mathcal{D}_{\xi}}\left[f(\mathbf{x};\xi)\right]$. For any $\xi\in\Xi$, the random loss function $f(\cdot;\xi)$ is $\alpha$-exp-concave, differentiable and $G$-Lipschitz over $\mathcal{X}$. For any query point $\mathbf{x}\in\mathcal{X}$, the stochastic gradient oracle returns $\nabla f(\mathbf{x};\xi)$ with $\xi$ i.i.d. drawn from $\mathcal{D}_{\xi}$.
[…]
Open problem [Koren, 2013].
Under Assumption˜4 and that $\mathcal{X}=\mathcal{B}(1)$, (a) Is it possible to find an SXO algorithm that attains the sample complexity of $\widetilde{O}(d/\epsilon)$ with only linear-in-$d$ runtime per iteration, i.e., $\widetilde{O}(d^{2}/\epsilon)$ runtime overall? (b) Is it possible to perform any better than $\widetilde{O}(d^{4}/\epsilon)$ runtime overall?
The first part of the open problem remains open. In particular, Mahdavi et al. [2015] establishes an information-theoretic sample complexity lower bound of $\Omega(d/\epsilon)$ for SXO, which implies a runtime lower bound of $\Omega(d^{2}/\epsilon)$ since each gradient query costs $\Omega(d)$ time.
Section˜4.2 answers the second part in the affirmative by combining LightONS with the online-to-batch conversion, obtaining total runtime $\widetilde{O}(d^{3}/\epsilon)$ while maintaining the optimal $\widetilde{O}(d/\epsilon)$ sample complexity. We further present evidence that this $\widetilde{O}(d^{3}/\epsilon)$ runtime is likely unimprovable. If confirmed, it would refute the first part and settle the full open problem.
4.2 Answering the Open Problem with LightONS
Built upon the online-to-batch conversion of Mehta [2017], the LightONS-based SXO method reduces the total runtime to $\widetilde{O}(d^{3}/\epsilon)$, with (near) optimal convergence rate both in high probability and in expectation, as the following theorem shows. Its proof is deferred to Section˜D.1.
Theorem 3.
Under Assumption˜4 and that $\mathcal{X}=\mathcal{B}(1)$, let LightONS (Algorithm˜3) run for $T$ rounds with gradients $\{\nabla f(\mathbf{x}_{t};\xi_{t})\}_{t=1}^{T}$ where $\{\mathbf{x}_{t}\}_{t=1}^{T}$ are decisions, let $\bar{\mathbf{x}}_{T}=\frac{1}{T}\sum_{t=1}^{T}\mathbf{x}_{t}$, then
  • For any $\delta\in(0,1)$, with probability at least $1-\delta$, when $T=\Theta\left(\frac{d}{\epsilon}\log\frac{d}{\epsilon}\log\frac{1}{\delta}\right)$,
    $F(\bar{\mathbf{x}}_{T})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\leq O(\epsilon)$ and $\textsc{Runtime}\leq\widetilde{O}\left(\frac{d^{3}}{\epsilon}\right)$.
  • When $T^{\prime}=\Theta\left(\frac{d}{\epsilon}\log\frac{d}{\epsilon}\right)$, $\mathop{\mathbb{E}}\left[F(\bar{\mathbf{x}}_{T^{\prime}})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\right]\leq O(\epsilon)$ and $\textsc{Runtime}\leq\widetilde{O}\left(\frac{d^{3}}{\epsilon}\right)$.
The LightONS-based SXO method in Theorem˜3 matches the best known SXO runtime up to poly-logarithmic factors. An online-to-batch conversion of OQNS also yields a total runtime of $\widetilde{O}(d^{3}/\epsilon)$ but carries significant practical overheads due to its large constant factors, as discussed in Section˜3. Our experiments in Appendix˜A, conducted under SXO settings, confirm this difference.
Conjecture on the optimality of $\widetilde{O}(d^{3}/\epsilon)$.
We conjecture that no SXO algorithm can asymptotically improve upon the total runtime of $\widetilde{O}(d^{3}/\epsilon)$. A natural alternative approach to SXO is empirical risk minimization (ERM) [Koren and Levy, 2015; Mehta, 2017], which reduces SXO to offline exp-concave optimization and admits offline acceleration techniques. Notably, even in the simplest least-squares setting with full access to the data, no $\widetilde{O}(d^{2}/\epsilon)$-time algorithm is known. A detailed discussion on offline acceleration techniques, including fast matrix multiplication and cutting-plane methods, is deferred to Section˜D.2. Moreover, ERM-based SXO methods require $\Omega(d^{2}+d/\epsilon)$ memory to store all samples, whereas the OXO-based methods require only $O(d^{2})$ memory.
Table 2: Regret and total runtime comparison between ONS and LightONS on three benchmark applications. “OQNS?” indicates whether OQNS supports the same application.
Application Regret (ONS) Runtime (ONS) Regret (LightONS) Runtime (LightONS) OQNS?
Gradient-norm adaptivity $O(d\log G_{T})$ $\widetilde{O}(d^{\omega}T)$ same $O(d^{2}T+d^{\omega}\sqrt{T}\log T)$ hardly
Generalized linear bandits $\widetilde{O}(d\sqrt{T}+\kappa d^{2})$ $\widetilde{O}(d^{2}KT\,+$ $d^{\omega}T)$ same order $\widetilde{O}(d^{2}KT\,+$ $d^{\omega}\cdot\min\{\sqrt{\kappa dT},T\})$ hardly
Memory-efficient OXO $O(d^{\prime}\log T)$ $\widetilde{O}(d^{\omega}T)$ same $O((d^{2}+d^{\prime}d\log T)T)$ intricate
5 Applications to Various Problems
ONS extends its influence far beyond the classical settings of OXO and SXO. Since LightONS is designed to retain core updates of ONS with minimal modifications, it preserves its elegant structure. This enables seamless integration into a range of applications where ONS serves as the computational core, not only maintaining its statistical advantages but also significantly enhancing efficiency.
We highlight three representative applications of ONS (gradient-norm adaptivity, parameteric stochastic bandits, and memory-efficient OXO) where LightONS seamlessly fits, preserving statistical benefits while requiring minimal additional analytical effort. In contrast, deploying OQNS in these settings is either infeasible or would require substantial and non-trivial modifications. The main results are summarized in Table˜2, comparing ONS with LightONS and OQNS. We defer further details and proofs to Appendix˜E.
5.1 Gradient-Norm Adaptivity
ONS naturally achieves a problem-dependent regret bound that scales with the accumulated squared gradient norms $G_{T}$ instead of the time horizon $T$:
$\textsc{Reg}_{T}(\mathbf{u})=O\left(d\log G_{T}\right),\quad\text{where}\penalty 10000\ G_{T}\triangleq\sum_{t=1}^{T}\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2},$ (12)
where $G_{T}$ can be $o(T)$ in benign environments and is upper bounded by $G^{2}T$, safeguarding the optimal regret $O(d\log T)$ against worst cases. Prior work has leveraged this gradient-norm adaptivity to achieve strong guarantees in various settings. For OXO with smooth loss functions, Orabona et al. [2012] achieve a small-loss adaptive regret bound $\textsc{Reg}_{T}=O(d\log L_{T})$, where $L_{T}\triangleq\min_{\mathbf{u}\in\mathcal{X}}\sum_{t=1}^{T}(f_{t}(\mathbf{u})-\min_{\mathbf{x}\in\mathcal{X}}f_{t}(\mathbf{x}))$. For OCO with unbounded domain, Cutkosky and Orabona [2018] achieve a comparator-norm adaptive regret bound without prior knowledge of the comparator norm $\left\|\mathbf{u}\right\|_{2}$, i.e., $\textsc{Reg}_{T}(\mathbf{u})=\widetilde{O}(\left\|\mathbf{u}\right\|_{2}\sqrt{dG_{T}})$ for any $\mathbf{u}\in\mathbb{R}^{d}$. Based on ONS, preceding results incur the worst-case runtime bottleneck $\widetilde{O}(d^{\omega}T)$. By incorporating LightONS, the runtime improves to $O(d^{2}T+d^{\omega}\sqrt{T}\log T)$ while preserving the same regret.
OQNS can hardly achieve gradient-norm adaptivity. Its computational efficiency relies on the log-barrier that stabilizes iterates for efficient matrix approximations. However, this log-barrier incurs an unavoidable $O(\log T)$ term which precludes adaptation to $G_{T}$. It is known that the OMD framework permits a flexible trade-off between stability and bias in the regret decomposition. While LightONS retains this flexibility, the log-barrier in OQNS overly suppresses the stability term, leading to a large bias term that does not scale with the gradient norms. Details are deferred to Section˜E.1.
5.2 Parametric Stochastic Bandits
Parametric stochastic bandits model decision-making problems with partial feedback, and generalized linear bandits (GLB) is a fundamental instance with non-linearly parameterized losses. A notable challenge in GLB is its dependence on the condition number $\kappa\propto\exp(D)$. Directly applying ONS to GLB yields an $\widetilde{O}(\kappa d\sqrt{T})$ regret [Jun et al., 2017]. Recently, Zhang et al. [2025] propose the first jointly efficient GLB algorithm, which achieves an $\widetilde{O}(d\sqrt{T}+\kappa d^{2})$ regret (statistically efficient) and is one-pass (computationally efficient). Their key technique is an ONS-based subroutine with non-monotonic local norms, which enjoys the one-pass efficiency of ONS and eliminates $\kappa$ from leading terms in regret. By replacing with a LightONS-based counterpart, we retain the same regret bound and improves the runtime from $\widetilde{O}((d^{2}K+d^{\omega})T)$ to $\widetilde{O}(d^{2}KT+d^{\omega}\cdot\min\{\sqrt{\kappa dT},T\})$, where $K$ is the number of arms for bandits. On condition that the regret is sublinear in $T$, i.e., $\kappa=o(T)$, our improvement yields a strictly lower runtime order.
OQNS can hardly be adapted to the method of Zhang et al. [2025]. The reasons are two-fold: (i) The GLB analysis treats the ONS-based subroutine as a white-box, exploiting negative terms in the OMD analysis, while OQNS fails to provide such analytical properties due to its deviation from OMD; (ii) The GLB method relies on non-monotonic local norms, while OQNS is tailored for monotonic local norms as in Eq.˜4, with which OQNS’s analysis is deeply coupled. On the contrary, the key ingredients of LightONS (deferred projection and domain conversion) are largely orthogonal to the core mirror-descent updates, seamlessly accommodating customized local norms and preserving necessary analytical properties for GLB. Details are deferred to Section˜E.2.
5.3 Memory-Efficient OXO
Another computational challenge of ONS is its $O(d^{2})$ working memory, in contrast to OGD’s $O(d)$ working memory. Luo et al. [2016] propose Sketched Online Newton Step (SON), which achieves linear-in-$d$ runtime and memory with matrix sketching when the gradient feedbacks exhibit low-dimensional structures. However, SON demands additional assumptions on the domain and loss functions and reverts to the $\widetilde{O}(d^{\omega}T)$ runtime under the standard OXO setting (Assumptions˜1, 2 and 3). Equipping LightONS with sketching, we propose a hybrid algorithm which combines the projection efficiency of LightONS with the memory efficiency of SON, extending the linear-in-$d$ runtime and memory to the standard OXO setting.
Although Mhammedi and Gatmiry [2023] suggest integrating sketching into OQNS, their algorithm’s considerable structural complexity may pose significant obstacles to such integration. Details are deferred to Section˜E.3.
6 Conclusion
For online exp-concave optimization (OXO), we propose LightONS, a simple yet powerful variant of ONS. By combining the deferred-projection mechanism with the improper-to-proper conversion, LightONS achieves significant reductions in worst-case total runtime while preserving the regret optimality of ONS. These gains extend to the stochastic optimization setting, answering a COLT’13 open problem on efficient and optimal SXO [Koren, 2013]. Moreover, due to its fidelity to the mirror-descent udpate of ONS, LightONS serves as an efficient drop-in replacement of ONS across diverse applications, including gradient-norm adaptivity, parametric stochastic bandits, and memory-efficient settings, all without compromising statistical guarantees.
Several important directions remain open. First, while LightONS adapts to gradient norms, achieving adaptivity to gradient variation for OXO within $\widetilde{O}(d^{2}T)$ runtime remains an open challenge [Chiang et al., 2012; Zhao et al., 2024]. Second, applying the LightONS technique to other preconditioned online learning algorithms may not improve their (asymptotic) runtime if the bottleneck lies in other operations, such as explicit matrix factorization rather than Mahalanobis projections (e.g., in AdaGrad [Duchi et al., 2011]). Accelerating these algorithms remains unclear. These future directions motivate the development of broader, general-purpose acceleration techniques for Hessian-related online learning algorithms.
Appendix A Empirical Validation
We conduct numerical experiments to validate the theoretical guarantees of LightONS, especially its negligible statistical gap from ONS and its non-asymptotic statistical advantages over OQNS.
Setup.
We evaluate on two fundamental tasks, linear and logistic regression, over the Euclidean ball domain $\mathcal{X}=\mathcal{B}(D/2)\subset\mathbb{R}^{d}$ over a time horizon of $T$. With $\{(\mathbf{x}_{t},y_{t})\}_{t=1}^{T}$ being i.i.d. sampled from the folded standard Gaussian distribution, i.e., each entry of $\mathbf{x}_{t}$ and $y_{t}$ is the absolute value of a standard Gaussian random variable, the (online) loss functions are designed as
$\ell^{\text{linear}}_{t}(\mathbf{w})=\frac{1}{2}\left(\sqrt{\frac{G}{D}}\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}+\frac{\sqrt{DG}}{2}y_{t}\right)^{2},\quad\ell^{\text{logistic}}_{t}(\mathbf{w})=\log\left(1+\exp\left(G\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}\right)\right).$
Both loss functions are $G$-Lipschitz and $\alpha$-exp-concave with $\alpha=1/(DG)$ and $\alpha=\exp(-DG)$, respectively. Our logistic regression setup slightly differs from the standard formulation, as we omit the binary label $y_{t}\in\{0,1\}$ for simplicity, which does not affect exp-concavity.
Three algorithms, LightONS, ONS, and OQNS, are configured with their theoretically optimal parameters. The implementation of ONS and LightONS follows Algorithms˜1 and 3, respectively; OQNS is implemented in accordance with Algorithm 3 of Mhammedi and Gatmiry [2023].
Results.
Figure˜1 displays all results averaged over 5 independent runs with the same seeds, where the averaged performance is shown as a dark line, while individual runs are in transparent lines.
Our experiments confirm that computational gains of LightONS incur negligible loss in statistical performance. As shown in Figures˜1(a), 1(d), 1(b) and 1(e), curves of LightONS (red dotted) overlap with those of ONS (green solid). This provides empirical evidence that LightONS retains the sharp regret constants of the ONS. In contrast, OQNS (blue dashed) shows noticeably higher regret, indicating that its theoretical bounds may carry greater constants.
The key advantage of LightONS, its computational efficiency, is evident in Figures˜1(c) and 1(f). After an initial convergence phase (e.g., the first $10^{2}$ rounds in Figure˜1(f)), both ONS and LightONS stabilize near the offline optimal solution. Then ONS frequently pushes decisions outside the domain, triggering costly Mahalanobis projections. In contrast, LightONS consistently outputs decisions with $\left\|\mathbf{y}_{t}\right\|\leq kD/2=2$ and avoids Mahalanobis projections.
Appendix B Technical Lemmas
In this section, we present technical lemmas used in the subsequent analysis.
B.1 Proof of Lemma˜1
Proof.
Since $g(\mathbf{x})=\mathrm{e}^{-\alpha f(\mathbf{x})}$ is concave and positive, we have that $h(\mathbf{x})=\mathrm{e}^{-2\gamma_{0}f(\mathbf{x})}$ is also concave. Specifically, for any $\lambda\in[0,1]$ and any $(\mathbf{x},\mathbf{y})\in\mathcal{X}^{2}$, it holds that
$\geq\lambda g(\mathbf{x})^{2\gamma_{0}/\alpha}+(1-\lambda)g(\mathbf{y})^{2\gamma_{0}/\alpha}=\lambda h(\mathbf{x})+(1-\lambda)h(\mathbf{y}),$
where the first inequality follows from the concavity of $g$ and the second from the concavity of $a\mapsto a^{2\gamma_{0}/\alpha}$ (where $2\gamma_{0}/\alpha\leq 1$). Next, by the concavity of $h$, for any $(\mathbf{x},\mathbf{u})\in\mathcal{X}^{2}$, we have
Rearranging the preceding inequality yields
To prove Eq.˜2, it suffices to apply that $\log(1+a)\leq a-a^{2}/4$ (which holds for any $\left|a\right|\leq 1$) to the right-hand side of the preceding inequality. We note that $\left|2\gamma_{0}\nabla f(\mathbf{x})^{\mathsf{T}}(\mathbf{x}-\mathbf{u})\right|\leq 1$ due to the selection of the curvature parameter $\gamma_{0}=\frac{1}{2}\min\{\frac{1}{DG},\alpha\}$. ∎
B.2 Elliptical Potential Lemmas
We introduce two lemmas tackling Hessian-related matrices, which are essential for both regret and runtime analysis. Lemma˜5 is standard in the ONS literature (e.g., [Hazan et al., 2007, Luo et al., 2016]). The proof of Lemma˜6 mirrors that of Lemma˜5, differing only in the potential function: $F(A)=-\operatorname{tr}(A)$ versus $F(A)=\log\operatorname{det}(A)$.
Lemma 5.
Lemma 6.
Proof of Lemmas˜5 and 6.
The first equality follows from $A_{i}-A_{i-1}=\mathbf{v}_{i}\mathbf{v}_{i}^{\mathsf{T}}$ (by the definition of $A_{i}$) and $\mathbf{u}^{\mathsf{T}}A\mathbf{v}=\left\langle A,\mathbf{v}\mathbf{u}^{\mathsf{T}}\right\rangle_{\operatorname{F}}$ (by the definition of the $\left\langle\cdot,\cdot\right\rangle_{\operatorname{F}}$); The second inequality from the concavity of $F$; The last equality from telescoping. Substituting the respective potentials into Eq.˜15 yields Lemmas˜5 and 6. Concretely,
  • For Lemma˜5, $F(A)=\log\operatorname{det}(A)$ and $\nabla{F}(A)=A^{-1}$.
  • For Lemma˜6, $F(A)=-\operatorname{tr}(A^{-1})$ and $\nabla{F}(A)=A^{-2}$.
@@ B.2 Elliptical Potential Lemmas
B.3 Numerical Lemmas
The overall efficiency of LightONS hinges on two key operations, namely the matrix inversion $A_{t}^{-1}$ and the (infrequent) Mahalanobis projection $\Pi^{A_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{y}}_{t+1}]$. In this subsection, we detail efficient numerical approaches for the two operations respectively.
Matrix inversion.
@@ B.3 Numerical Lemmas
In LightONS, all Mahalanobis projections are onto the Euclidean ball $\widetilde{\mathcal{X}}_{k}=\mathcal{B}(kD/2)$ rather than the potentially complex domain $\mathcal{X}$. This geometry enables customized numerical approaches faster than generic solvers [Lee et al., 2015, Jiang et al., 2020]. Specifically, the dual problem of the Mahalanobis projection reduces to a one-dimensional root-finding problem (Exercise 4.22 of Boyd and Vandenberghe [2004]), solvable via bisection.
@@
We propose FastProj in Algorithm˜4, which offers two implementation choices for different purposes. Choice 1 attains superior theoretical dependence on $d$ by exploiting fast matrix multiplication with exponent $\omega<2.3714$ [Alman et al., 2025]. Choice 2 offers stronger practical performance via tridiagonalization tricks [Parlett, 1998, Golub and Van Loan, 2013], which is more favorable in practice where $\omega=3$. The following lemma characterizes the error and runtime of FastProj.
Lemma 7.
Let $\mathbf{v}$ be the output of FastProj (Algorithm˜4), and $\mathbf{v}^{*}=\Pi^{A}_{\mathcal{B}(R)}[\mathbf{u}]$ be the exact Mahalanobis projection, then $\mathbf{v}\in\mathcal{B}(R)$ and $\left\|\mathbf{v}-\mathbf{v}^{*}\right\|_{2}\leq\zeta$. The number of bisection $n=O(\log(\frac{1}{\zeta}\frac{\penalty 10000\ \overline{\lambda}\penalty 10000\ }{\penalty 10000\ \underline{\lambda}\penalty 10000\ }))$. With choice 1, the runtime is $O(d^{\omega}n)$, and with choice 2, the runtime is $O(d^{3}+dn)$.
@@
In the context of LightONS (Algorithm˜3), it suffices to set the error tolerance as $\zeta_{t}=O(1/t^{2})$ at the $t$-th round. The accumulated truncation error contributes only a negligible additive constant to the regret of LightONS. A detailed discussion is deferred to Lemma˜12 in Section˜C.4. Moreover, $\underline{\lambda}_{t}=\epsilon$ and $\overline{\lambda}_{t}=\epsilon+c_{g}^{2}G^{2}t$, we have $n_{t}=O(\log t)$ at the $t$-th round.
To prove Lemma˜7, we first present two supporting lemmas. Lemma˜8 reduces the Mahalanobis projection onto an Euclidean ball to a one-dimensional root-finding problem. Lemma˜9 bridges the truncation error of root-finding and the truncation error of Mahalanobis projection.
Lemma 8.
$\rho(\mu)=\left\|\left(A+\mu I\right)^{-1}A\mathbf{y}\right\|_{2}^{2}-R^{2}=\sum_{i=1}^{d}\frac{v_{i}^{2}}{\left(1+\frac{\mu}{\lambda_{i}}\right)^{2}}-R^{2},$ (17)
and $\mu^{\ast}$ satisfies that
$\left(\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1\right)\lambda_{d}<\mu^{\ast}<\left(\frac{\left\|\mathbf{y}\right\|_{2}}{R}-1\right)\lambda_{1},$ (18)
where $v_{i}=\mathbf{e}_{i}^{\mathsf{T}}Q^{\mathsf{T}}\mathbf{y}$, $\mathbf{e}_{i}$ is the $i$-th standard basis vector, and $A=Q\Lambda Q^{\mathsf{T}}$ is the eigendecomposition, $\Lambda=\operatorname{diag}\left(\lambda_{1},\dots,\lambda_{d}\right)$ with $\lambda_{1}\geq\dots\geq\lambda_{d}>0$.
Proof of Lemma˜8.
The Mahalanobis projection onto an Euclidean ball is formulated as